# MIT-CTP/5723

### v2

@@ C.3 Generative modelling
The $t\bar{t}+n\;\mathrm{jets}$, $n=0...4$ dataset is simulated with the MadGraph 3.5.1 [3] event generation toolchain, consisting of MadEvent [3] for the underlying hard process, Pythia 8 [71] for the parton shower, Delphes 3 [34] for a fast detector simulation, and the anti-kT jet reconstruction algorithm [20] with $R=0.4$ as implemented in FastJet [21]. The Pythia simulation does not include multi-parton interactions. We use the ATLAS detector card for the Delphes detector simulation, apply the standard phase space cuts $p_{T}>20\;\mathrm{GeV},\Delta R=\sqrt{\Delta\phi^{2}+\Delta\eta^{2}}<0.4$ and require 2 $b$-tagged jets. The events are reconstructed with a $\chi^{2}$-based reconstruction algorithm [2], and identical particles are ordered by $p_{T}$.
The sizes of the $t\bar{t}+n\;\mathrm{jets}$, $n=0...4$ datasets reflect the frequency of the respective processes, resulting in $9.8\times 10^{6}$ ($n=0$), $7.2\times 10^{6}$ ($n=1$), $3.7\times 10^{6}$ ($n=2$), $1.5\times 10^{6}$ ($n=3$) and $4.8\times 10^{5}$ ($n=4$) events. On each dataset, 1% of the samples are set aside as validation and test split. We rescale the four-momenta $p$ by the standard deviation of all five datasets $206.6\;\mathrm{GeV}$ for processing with neural networks.
@@
The L-GATr score network operates in Minkowski space $p=(E,p_{x},p_{y},p_{z})$, whereas flow matching happens in the physically motivated coordinates $y=(y_{m},y_{p},\eta,\phi)$ defined in Eq. (4). After transforming $y$ into $p$, we embed each particle $p$ into geometric algebra representations. We use scalar channels for the one-hot-encoded particle type and the flow time, for which we use a sinusoidal embedding with 8 channels [72]. We add extra multivector channels as symmetry-breaking inputs, in particular the time direction $(1,0,0,0)$ as a rank-1 multivector, and the plane orthogonal to the beam direction encoded as a rank-2 multivector. The time direction is required to break the non-compact special orthochronous Lorentz group $SO^{+}(1,3)$ down to the compact subgroup $SO(3)$, as discussed in Sec. 3. The output of the L-GATr network is a vector field in Minkowski space $(v_{E},v_{p_{x}},v_{p_{y}},v_{p_{z}})$. To obtain vector fields in the manifold coordinates $y$, we multiply with the Jacobians of the transformation $p\to y$ to obtain $(v_{y_{m}},v_{y_{p}},v_{\eta},v_{\phi})$. Because of the logarithm in the change of variables, the Jacobians for $y_{m},y_{p}$ can take on large values, which makes training unstable. To avoid this complication, we extract $\tilde{v}_{y_{m}},\tilde{v}_{y_{p}}$ directly from the scalar output channels of L-GATr and use these values as vector field $v_{y_{m}},v_{y_{p}}$. Like the symmetry-breaking inputs, this procedure breaks the Lorentz symmetry down to the residual symmetry group of the measurement process. We find it beneficial to perform the transformation $y\leftrightarrow p$ at 64-bit floating-point precision, which has no noticable effect on the computational cost. We use an L-GATr network with 16 multivector channels, 32 scalar channels, 6 L-GATr blocks, and 8 attention heads, totalling $5.4\times 10^{5}$ learnable parameters.
The Transformer score network directly operates in the physically motivated $y$ space. Each particle is embedded into one token, with 4 channels for the components of $y$, 8 channels for the flow time in a sinusoidal embedding [72], and the one-hot-encoded particle type. The network has 108 channels, 6 Transformer blocks, and 8 attention heads, totalling to $5.7\times 10^{5}$ learnable parameters.
@@
All networks are trained for $2\times 10^{5}$ iterations with batchsize 2048 using the Adam optimizer with default settings and an initial learning rate of $0.001$. We evaluate the validation loss every $10^{3}$ iterations and decrease the learning rate by a factor of 10 after no improvements for 20 validation steps. We stop training after the validation loss has not improved after 50 validation steps, and pick the best network with the best validation loss.
@@
The base distribution is defined in the rescaled Minkowski space discussed above. We use unit Gaussians for the spatial momentum $p_{x,y,z}\sim\mathcal{N}(0,1)$ and the log-transformed squared mass $y_{m}=\log m^{2}\sim\mathcal{N}(0,1)$. We ensure the constraints $p_{T}>20\;\mathrm{GeV},\Delta R>0.4$ through rejection sampling. We have experimented with other base distributions and find similar performance.
@@
$\frac{d}{dt}\begin{pmatrix}x_{t}\\ f_{t}(x)\end{pmatrix}=\begin{pmatrix}v_{t}(x)\\ -\mathrm{div}(v_{t})(x_{t})\end{pmatrix}$ (9)
@@
We train a MLP classifier to distinguish generated events from the ground truth. The classifier inputs are full events in the $y$ representation, together with challenging correlations, in particular all pairwise $\Delta R$ values, and the $y$ representations of the reconstructed intermediate particles $t,\bar{t},W^{+},W^{-}$. The classifier network has 256 channels and 3 layers. It is trained for 500 epochs with batchsize 1024, a dropout rate of 0.1, and the Adam optimizer with default settings. We start with an initial learning rate of 0.0003 and decrease the learning rate after no improvements in the validation loss for 5 epochs. We stop training after 10 epochs without improvements in the validation loss and load the best-validation model afterwards. We use the full ground truth dataset as well as 1M generated events and split into 80% for training and 10% each for testing and validation. We use the AUC of this classifier evaluated on the test dataset as a scalar metric, with the value $0.5$ for a perfect generator. Neural classifiers approximate the event-wise likelihood ratio $p_{\mathrm{data}}(x)/p_{\mathrm{model}}(x)$ of single events, which is the most powerful test statistic according to the Neyman-Pearson lemma, and opens many ways to further study the performance of the generator beyond scalar metrics [33].
@@ C.4 Computational cost and scalability
Figure 7: Inference cost (wall-time per forward pass) as a function of the number of particles. We compare L-GATr, a Transformer, and a message-passing graph neural network (we use CGENN [67] but expect similar results for other architectures). The latter runs out of memory when evaluating more than a few hundred particles. While we do our best to find comparable settings, such comparisons depend on a lot of choices and should be interpreted with care. Nevertheless, we believe they illustrate that L-GATr scales to large systems like a Transformer, thanks to it being based on dot-product attention.
In Fig. 7 we compare L-GATr to a message-passing graph neural network (we use CGENN [67]) and a vanilla Transformer in terms of their test-time computational costs. For this comparison, we use small versions of all architectures, consisting of a single model block and around $9\times 10^{4}$ learnable parameters. In the case of the Transformer and L-GATr, we fix their layer structure so that inputs going into the attention layer consist of 72 channels. Our measurements are performed with datasets made up by a single sample and all models are run on an H100 GPU.
@@
Compared to an equivariant graph network, L-GATr is clearly more efficient in terms of compute time and memory. Already at small systems, L-GATr can be evaluated an order of magnitude faster. The difference is even more pronounced in terms of memory: the graph network ran out of memory for more than a few hundred particles. This is largely because graph network implementations are often optimized for sparse computational graphs, but here we use fully connected graphs: LHC problems often benefit from a fully connected computational graph, because pairwise interactions do not usually decay with the Minkowski norm of the distance. Transformer-based approaches like L-GATr can thus have substantial computational advantages in particle physics problems that involve a large number of particles.

### v3

@@ C.3 Generative modelling
The $t\bar{t}+n\;\mathrm{jets}$, $n=0...4$ dataset is simulated with the MadGraph 3.5.1 event generation toolchain, consisting of MadEvent [4] for the underlying hard process, Pythia 8 [72] for the parton shower, Delphes 3 [35] for a fast detector simulation, and the anti-kT jet reconstruction algorithm [21] with $R=0.4$ as implemented in FastJet [22]. The Pythia simulation does not include multi-parton interactions. We use the ATLAS detector card for the Delphes detector simulation, apply the phase space cuts $p_{T}>22\;\mathrm{GeV},|\eta|<5,\Delta R=\sqrt{\Delta\phi^{2}+\Delta\eta^{2}}<0.5$ and require 2 $b$-tagged jets. The events are reconstructed with a $\chi^{2}$-based reconstruction algorithm [2], and identical particles are ordered by $p_{T}$.
The sizes of the $t\bar{t}+n\;\mathrm{jets}$, $n=0...4$ datasets reflect the frequency of the respective processes, resulting in $9.8\times 10^{6}$ ($n=0$), $7.2\times 10^{6}$ ($n=1$), $3.7\times 10^{6}$ ($n=2$), $1.5\times 10^{6}$ ($n=3$) and $4.8\times 10^{5}$ ($n=4$) events. On each dataset, 1% of the events are set aside as validation and test split. We rescale the four-momenta $p$ by the standard deviation of all five datasets $206.6\;\mathrm{GeV}$ for processing with neural networks.
@@
The L-GATr score network operates in Minkowski space $p=(E,p_{x},p_{y},p_{z})$, whereas flow matching happens in the physically motivated coordinates $y=(y_{m},y_{p},\eta,\phi)$ defined in Eq. (4). After transforming $y$ into $p$, we embed each particle $p$ into geometric algebra representations. We use scalar channels for the one-hot-encoded particle type and the flow time, for which we use a Gaussian Fourier Projection with 8 channels [73]. We add the same symmetry breaking inputs that we used for the top tagging task, this time as extra multivector channels. The output of the L-GATr network is a vector field in Minkowski space $(v_{E},v_{p_{x}},v_{p_{y}},v_{p_{z}})$. To obtain vector fields in the manifold coordinates $y$, we multiply with the Jacobians of the transformation $p\to y$ to obtain $(v_{y_{m}},v_{y_{p}},v_{\eta},v_{\phi})$. Because of the logarithm in the change of variables, the Jacobians for $y_{m},y_{p}$ can take on large values, which makes training unstable. To avoid this complication, we extract $\tilde{v}_{y_{m}},\tilde{v}_{y_{p}}$ directly from the scalar output channels of L-GATr and use these values as vector field $v_{y_{m}},v_{y_{p}}$. Like the symmetry-breaking inputs, this procedure breaks the Lorentz symmetry down to the residual symmetry group of the measurement process. We find it beneficial to perform the transformation $y\leftrightarrow p$ at 64-bit floating-point precision, which has no noticable effect on the computational cost. We use an L-GATr network with 16 multivector channels, 32 scalar channels, 6 L-GATr blocks, and 8 attention heads, totalling $5.4\times 10^{5}$ learnable parameters.
We implement the E(3)-GATr score network following the same prescription outlined above for L-GATr. The only difference is that we only transform between $y$ and $(y_{m},p_{x},p_{y},p_{z})$ to avoid the large jacobians from the $y_{m}\leftrightarrow m$ transformation. Our task only requires SO(3)-equivariance, therefore we do not use of the translation-equivariant input and output representations of E(3)-GATr, but we keep the internal translation-equivariant representations. For each particle, we embed the three-momentum $(p_{x},p_{y},p_{z})$ into the vector component of the multivector, and $y_{m}$ as a scalar.
The Transformer score network directly operates in the physically motivated $y$ space. Each particle is embedded into one token, with 4 channels for the components of $y$, 8 channels for the flow time in a Gaussian Fourier Projection, and the one-hot-encoded particle type. The network has 108 channels, 6 Transformer blocks, and 8 attention heads, totalling to $5.7\times 10^{5}$ learnable parameters.
@@
We implement the JetGPT model closely following Ref. [18]. JetGPT is an autoregressive mixture model, i.e. its parameters are predicted autoregressively for each component of the high-dimensional distribution. We use the same transformer as for the Transformer-CFM to predict the mixture parameters, but using a triangular attention matrix to achieve the autoregressive structure. The components are the same $y$ coordinates that we use for the flow matching models, amounting to four tokens for each particle. We use gaussian mixture models for the non-periodic coordinates $(y_{m},y_{p},\eta)$, and von Mises mixture models for the periodic coordinate $\phi$. Due to the autoregressive structure, the ordering of components affects the performance. As more conditions are added, the components will be harder to learn using a fixed amount of training data. We order the particles as $(t_{1},q_{1},q_{2},t_{2},q_{3},q_{4},j_{1},\dots j_{4})$, and within each particle we order the components as $(\phi,y_{p},\eta,y_{m})$. We train the model to minimize the joint log-likelihood and sample the components sequentially. Each component is embedded into one token, with 1 channel for the value, and 24 to 40 channels for the one-hot-encoded component type. We use the same transformer as for the Transformer-CFM, with 108 channels, 6 Transformer blocks, and 8 attention heads, but the last layer maps onto the 108 parameters of a mixture model of 36 gaussian or von Mises distributions.
@@
All networks are trained for $2\times 10^{5}$ iterations with batchsize 2048 using the Adam optimizer with default settings and an initial learning rate of $0.001$. We evaluate the validation loss every $10^{3}$ iterations and decrease the learning rate by a factor of 10 after no improvements for 20 validation steps. We stop training after the validation loss has not improved after 50 validation steps, and pick the network with the best validation loss.
@@
The base distribution is defined in the rescaled Minkowski space discussed above. We use standardized Gaussians for the spatial momentum $p_{x,y,z}\sim\mathcal{N}(0,1)$ and the log-transformed squared mass $y_{m}=\log m^{2}\sim\mathcal{N}(0,1)$. We ensure the constraints $p_{T}>22\;\mathrm{GeV},\Delta R>0.5$ through rejection sampling. We have experimented with other base distributions and find similar performance.
@@
$\frac{d}{dt}\begin{pmatrix}x_{t}\\ f_{t}(x_{t})\end{pmatrix}=\begin{pmatrix}v_{t}(x_{t})\\ -\mathrm{div}(v_{t})(x_{t})\end{pmatrix}$ (9)
@@
We train a MLP classifier to distinguish generated events from the ground truth. The classifier inputs are full events in the $y$ representation, together with challenging correlations, in particular all pairwise $\Delta R$ values, and the $y$ representations of the reconstructed particles $t,\bar{t},W^{+},W^{-}$. The classifier network has 256 channels and 3 layers. It is trained for 500 epochs with batchsize 1024, a dropout rate of 0.1, and the Adam optimizer with default settings. We start with an initial learning rate of 0.0003 and decrease the learning rate by a factor of 10 after no improvements in the validation loss for 5 epochs. We stop training after 10 epochs without improvements in the validation loss and load the best-validation model afterwards. We use the full ground truth dataset as well as 1M generated events and split into 80% for training and 10% each for testing and validation. We use the AUC of this classifier evaluated on the test dataset as a scalar metric, with the value $0.5$ for a perfect generator. Neural classifiers approximate the event-wise likelihood ratio $p_{\mathrm{data}}(x)/p_{\mathrm{model}}(x)$ of single events, which is the most powerful test statistic according to the Neyman-Pearson lemma, and opens many ways to further study the performance of the generator beyond scalar metrics [34].
@@ C.4 Computational cost and scalability
Figure 7: Inference cost (wall-time per forward pass) as a function of the number of particles. We compare L-GATr, a Transformer, and a message-passing graph neural network (we use CGENN [68] but expect similar results for other architectures). The latter runs out of memory when evaluating more than a thousand particles. While we do our best to find comparable settings, such comparisons depend on a lot of choices and should be interpreted with care. Nevertheless, we believe they illustrate that L-GATr scales to large systems like a Transformer, thanks to it being based on dot-product attention.
In Fig. 7 we compare L-GATr to a message-passing graph neural network (we use CGENN [68]) and a vanilla Transformer in terms of their test-time computational costs. For this comparison, we use small versions of all architectures, consisting of a single model block and around $2\times 10^{5}$ learnable parameters. In the case of the Transformer and L-GATr, we fix their layer structure so that inputs going into the attention layer consist of 72 channels. Our measurements are performed with datasets made up by a single sample and all models are run on an H100 GPU.
@@
Compared to an equivariant graph network, L-GATr is clearly more efficient in terms of compute time and memory. Already at small systems, L-GATr can be evaluated an order of magnitude faster. The difference is even more pronounced in terms of memory: the graph network ran out of memory for more than a thousand particles. This is largely because graph network implementations are often optimized for sparse computational graphs, but here we use fully connected graphs: LHC problems often benefit from a fully connected computational graph, because pairwise interactions do not usually decay with the Minkowski norm of the distance between momenta. Transformer-based approaches like L-GATr can thus have substantial computational advantages in particle physics problems that involve a large number of particles.